OpenStudy (anonymous):

If you can point me in the right direction, that would be wonderful! If f(x)=8sin^-1(x^3) (the inverse of sin) What is f'(x)?

OpenStudy (richyw):

\[8\sin^{-1}(x^3)\]

OpenStudy (richyw):

do you know what the derivative of \(\arcsin(x)\) is?

OpenStudy (anonymous):

I'm not sure...I know that you look at the right triangle it forms and that the sides are x=x^3, the hyp.=1 and y=sqrt(1-x^2) ..